Solve the IVP (for the Bernoulli equation):

dy/dx − (1/x)y = 1/y , y(1) = −3
Answer #1

ou e » y dx - y = Let y2 = 2 - 2y d y = d7 Y10)=-3 - y d y = 2 2 2 2 I from differential equation is put >> - 1 - 22 - 27 = 2I = C-2 = G = 11 So, Initial value: [Y(0= - J 11.22-22]